OpenStudy (anonymous):

4(t^-5) Find f'(t). Find f'(1). Help would be great!

OpenStudy (anonymous):

f'(t) = -20t^(-6) So, f'(1) = -20(1)^(-6) = - 20.

OpenStudy (anonymous):

Use basic differentiation rules to solve this. Use the power rule to find that the derivative of (t^-5) is -5t^-6. This is then multiplied by -4 and rewritten as -20/t^6. f'(1) would therefore be -20.